OpenStudy (anonymous):

given f(2)=1 and the fact that g(x)=-2f(x-1)+5 to find g(3)

OpenStudy (amistre64):

when x=3, we know f(3-1) ....

OpenStudy (anonymous):

Sorry I don't understand?

OpenStudy (amistre64):

the information provided is rather simple to follow ... f(2) = 1 what is f(3-1) ?

OpenStudy (anonymous):

Where did you get f(3-1) ?

OpenStudy (anonymous):

Sorry I don't understand?

OpenStudy (amistre64):

g(x) relates to g(3) by letting x=3

OpenStudy (amistre64):

since they give you: \[g(x)=2~\color{red}{f(x-1)}+5\] and say to find g(\(\color{orange}3\))

OpenStudy (amistre64):

\[g(x)=-2~\color{red}{f(x-1)}+5\] \[g(3)=-2~\color{red}{f(3-1)}+5\] \[g(3)=-2~\color{red}{f(2)}+5\]

OpenStudy (amistre64):

they say f(2) = 1 sooo .... \[g(3)=-2~\color{red}{f(2)}+5\] \[g(3)=-2~\color{red}{*1}+5\] \[g(3)=-2+5\]

OpenStudy (anonymous):

Ok thank you!

OpenStudy (amistre64):

yep